# Reinforced Prompt Personalization for Recommendation with Large Language Models

### v1

@@ 4.4. Generalization ability (RQ2)
To answer RQ2, we extend the evaluation of RPP/RPP+ on diverse frozen LLMs (i.e., LLaMa2-7B-chat, ChatGPT, and Alpaca) and compare the performance of RPP/RPP+ with task-wise prompt-based baselines. The results are listed in Table 3. RPP/RPP+ can generalize well on diverse LLMs and prominently enhance the ranking performance of LLMs. For example, RPP outperforms enumeration in terms of the NDCG@10 metric on LLaMa-7B-chat, ChatGPT, and fine-tuned Alpaca by $0.48$, $0.24$, and $0.53$, respectively, on the Lastfm dataset. The relatively less significant impact on ChatGPT’s performance might be attributed to its stronger robustness to prompts compared with other LLMs. The consistent performance improvement across the three types of frozen LLMs demonstrates the effectiveness and generalization ability of RPP/RPP+.
The results for Q1 and Q2 underscore the importance of instance-wise prompting and the effectiveness of RPP/RPP+ from a practical standpoint, confirming our viewpoint that personalizing instance-wise prompts can enhance LLMs’ recommendation ability. Furthermore, dynamically refining actions during the iterative process can lead to a slight performance improvement of RPP+ compared to RPP, possibly due to more flexible and higher-quality prompts after refinement. However, the improvement is not significant and will consume additional computational resources. Consequently, the trade-off between efficiency and performance can be considered when choosing between the RPP and RPP+ in practice.
@@
Figure 3. The performance comparison between task-wise and instance-wise prompt-based methods with different patterns. “Manual” and “Enum” represent the task-wise prompt-based methods. “RPP-Ro”, “RPP-Hi”, “RPP-Re”, and “RPP-Ou” are the 4 variations of RPP/RPP+ on “role-playing”, “history records”, “reasoning guidance”, and “output format” patterns, respectively, representing instance-wise prompt-based methods.
@@ 4.5. Ablation Study (RQ3)
The ablation study explores the impact of personalizing each pattern (i.e., “role-playing”, “history records”, “reasoning guidance,” and “output format”) on facilitating LLMs’ recommendation ability. Specifically, we perform experiments on the frozen LLaMa2-7B-chat with 4 variations of RPP/RPP+, where only one pattern in the prompts is personalized by the corresponding agent. As shown in Table 4,  5, and Figure 3, personalizing any of the patterns can enhance the performance of LLMs to rank items, and the full framework of RPP/RPP+ attains the highest performance. Personalizing “reasoning guidance” yields the greatest benefits compared with other patterns, providing interpretability for recommendations and demonstrating the effectiveness of the personalized reasoning process in accurately analyzing user preferences. Besides, “role-playing” and “output format” underscore LLMs’ inherent sensitivity to prompts that modifications to the prompt expression can lead to LLMs’ better output. The “history records” confirms the impact of users’ short or long-term interaction history on recommendations. This emphasizes that decomposing the prompt into four patterns and individually tailoring them can contribute to LLMs’ recommendations.
@@ 4.6. Sensitivity Analysis (RQ4)
We investigate the effect of the number of training examples (users) on RPP/RPP+ with LLaMa2-7B-chat in Figure 4. As the number of training examples increases from $100$ to $500$, the performance of RPP/RPP+ initially improves and subsequently stabilizes. Considering both the performance and resource consumption, we set the number of training examples as 200. Additionally, we explore the sensitivity of RPP/RPP+ to the number of candidate items within the range of $\{5,10,15,20\}$ in Figure 5. Observing the curves corresponding to the enumeration method and our framework RPP/RPP+, the ranking performance of LLMs drops as the number of candidate items increases, which can be attributed to the limitation of LLMs in processing long inputs. However, the capability of RPP/RPP+ to enhance the recommendation performance of LLMs is robust since the gap between the curves of RPP/RPP+ and enumeration remains stable. This observation highlights the effectiveness and stability of personalizing instance-wise prompts rather than solely relying on task-wise prompts.
@@ 4.7. Case Study
Through the case study presented in Figure 678, and 9, we provide an intuitive demonstration of how variations in these prompts impact the ranking results. For the user in Figure 6, personalizing the “role-playing” pattern from “You are a movie expert” to “ You are a movie ranker to match user’s interest” results in a higher ranking of the ground truth item that the user interacts with next. This illustrates that assigning LLMs as movie rankers to analyze users’ preferences can lead to more precise recommendations for this user. By personalizing the ’history record’ pattern from 5 interaction records to 2 historical records, the ranking results become more precise for the user in Figure 7. This demonstrates that considering a shorter-term interaction history is more effective in capturing his preferences than a long-term one. For the user in Figure 8, personalizing the “reasoning guidance” pattern from “think step by step” to “ update the film preferences before ranking” leads to better recommendations, proving that utilizing LLMs’ refinement and update mechanism is more reasonable than deducing without specific instructions for him. Moreover, personalizing the “output format” pattern from “giving the output format example” to “answering without any unnecessary steps” makes it easier for LLMs to understand the instructions and recommend for the user in Figure 9. We can observe that due to LLMs’ sensitivity to prompts and users’ inner difference, personalizing prompts for each user enables LLMs to generate recommendation results that better satisfy the users, with the ground truth prioritized. Furthermore, we visualize the distribution of users across different patterns under our experiments’ setting to illustrate user diversity vividly, as depicted in Figure 101112, and 13, underscoring the necessity of personalizing each pattern in the prompts for different users. This further validates the significance of instance-wise prompting for recommendation tasks and the effectiveness of our framework.
@@ 4.8. Timing and Computational Complexity Analysis
Compared to directly applying fixed prompts, the streamlined architecture of MARL, with parameters negligible relative to LLMs, incurs minimal additional computational and time costs. When customizing the prompt for each user, compared with heuristical crafting methods like enumeration or token-level optimization approaches like RLPrompt (Deng et al., 2022) and AutoPrompt (Shin et al., 2020), the sentence-level optimization can improve searching efficiency and obtain optimal prompts with minimal iterations, thus reducing timing and computational complexity. However, the multi-round iteration with LLMs is still required since achieving the desired prompt in a single step with MARL is challenging. Formally, let $N$ denote the number of users, $L$ denote the length of a prompt, $D$ denote the possible tokens, and $S$ demote the limited actions. RPP/RPP+ reduces the timing and computational complexity of designing prompts from $O(NLD)$ to $O(NS)$. Our experiments reveal that ChatGPT can complete prompt optimization in approximately three rounds, whereas LLaMa2-7B-chat requires two, seven, and eight iterations on three distinct datasets (i.e., ML-1M, Games, and Lastfm), and Alpaca takes around 8 iterations. The timing and computational complexity are acceptable to achieve significant performance improvements. It reveals that RPP and RPP+ can achieve a balance between prompt quality and resource efficiency.
@@ 5.1. LLMs for Recommender Systems
Recently, large language models (LLMs) have demonstrated remarkable capabilities in instruction following and intent reasoning, drawing researchers’ attention to combining LLMs with recommendations. The roles of LLMs at different stages of the recommendation pipeline can be classified into four categories (Lin et al., 2023): feature engineering (Borisov et al., 2023), feature encoder (Hou et al., 2023), scoring function (Li et al., 2023a), and pipeline controller (Gao et al., 2023). In our study, we conceptualize LLMs as RSs, wherein the LLMs directly generate recommendation outcomes in the role of scoring function. This mechanism relies heavily on the quality of the prompts designed. Tailoring prompts of LLMs for the recommendation tasks has been explored by recent research. InstructRec (Zhang et al., 2023b) develops a general instruction format that contains users’ preferences, intentions, task requirements, and contextual information. Chat-REC (Gao et al., 2023) recasts user profiles and prior interactions into prompts, enhancing the interactivity and explainability of the recommendation process. LLM4RS (Dai et al., 2023) designs domain-specific prompt format to analyze ChatGPT’s recommendation ability on all three ranking policies, including point-wise, pair-wise, and list-wise ranking. An evaluation study (Kang et al., 2023) investigates the recommendation ability of various LLMs by designing zero and few shots LLM prompts for rating prediction. P5 (Geng et al., 2022) converts interaction history, item metadata, and user review into a common format of prompt to realize various recommendation tasks. Most of them pay attention to developing task-wise prompts that are shared by all users in the specific task. Instead of designing a common prompt format for the specific recommendation task (Geng et al., 2022; Zhang et al., 2023b), we personalize instance-wise prompt for each user to explore the diversity of users’ preferences and intents.
@@ 5.2. Prompt Sensitivity of LLMs
Recent studies (Lu et al., 2022; Zhu et al., 2023) have examined the sensitivity of LLMs to prompts that even minor variations in prompts can lead to significantly different outputs, indicating that the one-size-fits-all prompt may not be ideal for different inputs. For example, KATE (Liu et al., 2022) confirmed that the number and order of in-context examples can influence the output of LLMs. Besides, variations in prompt formats within task descriptions may result in LLMs interpreting input instances differently, consequently influencing the outcomes (Zhao et al., 2021). These studies underscore the importance of optimizing prompts for specific inputs to enhance the performance of LLMs. Fixed prompt templates that assume uniformity across all samples in a task, regardless of their varying difficulty, may impair LLMs’ abilities due to a lack of instance-specific knowledge (Jiang et al., 2022). To address the difference in inputs, recent research has proposed instance-level prompt optimization (Jin et al., 2023; Wu et al., 2022), which involves rewriting the prompt for each input to better leverage the LLM’s capabilities for specific instances. IDPG (Wu et al., 2022) introduced a lightweight and trainable component to generate prompts based on each input instance. IPL (Jin et al., 2023) assumed that each learnable prompt token contributes differently to various instances, which can be learned by calculating relevance scores. To meet the specific demands for personalization in recommendation tasks, we enhance LLMs’ personalized recommendations by leveraging instance-wise prompts and introducing a method to optimize these prompts using MARL.
@@ 5.3. Discrete Prompt Engineering
Several approaches have explored prompt engineering in continuous and discrete space to obtain better answers from LLMs. Since the prompts for recent LLMs (ChatGPT, LLaMa2) are discrete and hard prompts, our work is strongly related to the discrete prompts engineering. Prior works have attempted to construct discrete prompts manually (Petroni et al., 2019; Schick and Schütze, 2021) or by heuristic (Jiang et al., 2020; Pan et al., 2023), which are limited by human effort or resource consumption. Recent research has explored constructing prompts automatically. For instance, prompt retrieval (Rubin et al., 2022) leverages annotated data and a language model, training a compact retriever to retrieve prompts based on input-output pairs. AutoPrompt (Shin et al., 2020) utilizes gradient-guided search to identify the optimal tokens within the prompt, although these prompts are typically not human-interpretable. RLPrompt (Deng et al., 2022) introduces a framework based on RL to generate prompts iteratively word-by-word from the vast vocabulary. An important distinction between our method and the existing methods is that we decompose the completed prompt into multiple key patterns, optimizing each pattern and concatenating them, which can enhance the efficiency of optimization and ensure the quality of prompts. To the best of our knowledge, our work is the first to explore instance-wise prompting in the domain of RSs.
@@ 7. Conclusion
In this paper, we propose instance-wise prompting to personalize prompts for individual users in recommendation tasks. To solve this task, we introduce a new framework, Reinforced Prompt Personalization (RPP/RPP+), leveraging multi-agent reinforcement learning to personalize essential patterns respectively and then concatenate them into a personalized prompt. We optimize the prompts at the sentence level and carefully design the expressions for each pattern, enhancing search efficiency and ensuring prompt quality. Experimental results validate the superiority of RPP/RPP+ over several traditional recommender models, few-shot methods, and task-wise prompting in ranking tasks, demonstrating the potential of personalizing instance-wise prompts.

### v2

@@ 4.4. Generalization ability (RQ2)
To answer RQ2, we extend the evaluation of RPP/RPP+ on diverse frozen LLMs (i.e., LLaMa2-7B-chat, ChatGPT, and Alpaca) and compare the performance of RPP/RPP+ with other prompt-based baselines. The results are listed in Table 4. RPP/RPP+ can generalize well on diverse LLMs and prominently enhance the ranking performance of LLMs. For example, RPP outperforms enumeration in terms of the NDCG@10 metric on LLaMa-7B-chat, ChatGPT, and fine-tuned Alpaca by $0.48$, $0.24$, and $0.53$, respectively, on the Lastfm dataset. The relatively less significant impact on ChatGPT’s performance might be attributed to its stronger robustness to prompts compared with other LLMs. The consistent performance improvement across the three types of frozen LLMs demonstrates the effectiveness and generalization ability of RPP/RPP+.
The results for Q1 and Q2 underscore the importance of instance-wise prompting and the effectiveness of RPP/RPP+ from a practical standpoint, confirming our viewpoint that personalizing instance-wise prompts can enhance LLMs’ recommendation ability. Furthermore, dynamically refining actions during the iterative process can lead to a slight performance improvement of RPP+ compared to RPP, possibly due to more flexible and higher-quality prompts after refinement. However, the improvement is not particularly significant ranging from 0.00 to 0.06 and requires additional computational resources during the LLM refinement process, the trade-off between efficiency and performance can be considered when choosing between the RPP and RPP+ in practice.
@@
Figure 3. The performance comparison between prompt-based methods and RPP/RPP+ with different patterns. “Manual”, “Enum”, and “GRIPS” represent the baseline prompt-based methods. “RPP-Ro”, “RPP-Hi”, “RPP-Re”, and “RPP-Ou” are the 4 variations of RPP/RPP+ on “role-playing”, “history records”, “reasoning guidance”, and “output format” patterns, respectively.
@@
(a) Ablation study for multi-agent reinforcement learning.
(b) Sensitivity to the number of sentence actions for each pattern.
Figure 6. Figure (a) is the performance comparison between multi-agent reinforcement learning and a single PPO for personalizing prompts. Figure (b) demonstrates RPP/RPP+’s sensitivity to the number of actions under each pattern.
@@ 4.5. Ablation Study (RQ3)
The ablation study explores the impact of personalizing each pattern (i.e., “role-playing”, “history records”, “reasoning guidance,” and “output format”) on facilitating LLMs’ recommendation ability. Specifically, we perform experiments on the frozen LLaMa2-7B-chat with 4 variations of RPP/RPP+, where only one pattern in the prompts is personalized by the corresponding agent. As shown in Table 5,  6, and Figure 3, personalizing any of the patterns can enhance the performance of LLMs to rank items, and the full framework of RPP/RPP+ attains the highest performance. Personalizing “reasoning guidance” yields the greatest benefits compared with other patterns, providing interpretability for recommendations and demonstrating the effectiveness of the personalized reasoning process in accurately analyzing user preferences. Besides, “role-playing” and “output format” underscore LLMs’ inherent sensitivity to prompts that modifications to the prompt expression can lead to LLMs’ better output. The “history records” confirms the impact of users’ short or long-term interaction history on recommendations. This emphasizes that decomposing the prompt into four patterns and individually tailoring them can contribute to LLMs’ recommendations. Furthermore, as some recommendation tasks focus on user modeling, we define a new pattern called "user profile" and conduct experiments to evaluate the effectiveness of RPP/RPP+ in personalizing this pattern using the ML-1M dataset. We utilize users’ profile features, such as gender, age, and occupation, as potential actions for the agent, which can choose whether to include each feature in the prompts. As shown in Table 7, personalizing the user profile leads to performance improvements compared with prompt-based baselines, highlighting the effectiveness of designating "user profile" as a distinct pattern and the advantages of personalizing prompts with RPP/RPP+.
To validate the design of “GRU” in Equation 4, we conduct ablation experiments to replace “GRU” as “mean pooling”. As presented in table 8, the performance of “RPP-pool” is worse than that of “RPP-Bert”, which validates the effectiveness of the “GRU” module since it can catch the sequential information of ranking results. Additionally, to evaluate the effectiveness of MARL which can personalize different patterns with different expertise from different agents, we utilize a single agent which is designed as PPO. The results are demonstrated in Figure 6(a), where the PPO-1 uses a single PPO to personalize the entire prompts without dividing them into four patterns, while PPO-4 personalizes the four patterns uniformly with one PPO. The superior performance of RPP compared to PPO-1 and PPO-4 highlights the need for multiple agents to personalize each of the four patterns respectively.
@@ 4.6. Sensitivity Analysis (RQ4)
We investigate the effect of the number of training examples (users) on RPP/RPP+ with LLaMa2-7B-chat in Figure 4. As the number of training examples increases from $100$ to $500$, the performance of RPP/RPP+ initially improves and subsequently stabilizes. Considering both the performance and resource consumption, we set the number of training examples as 200. Additionally, we explore the sensitivity of RPP/RPP+ to the number of candidate items within the range of $\{5,10,15,20\}$ in Figure 5. Observing the curves corresponding to the task-wise prompting method (e.g., enumeration) and our framework RPP/RPP+, the ranking performance of LLMs drops as the number of candidate items increases, which can be attributed to the limitation of LLMs in processing long inputs. However, the capability of RPP/RPP+ to enhance the recommendation performance of LLMs is robust since the gap between the curves of RPP/RPP+ and Enumeration remains stable. This observation highlights the effectiveness and stability of personalizing instance-wise prompts rather than solely relying on task-wise prompts.
Furthermore, to evaluate the impacts of different numbers of sentences under each pattern, we conduct experiments for the four patterns with results presented in Figure 6(b). The curves rise initially and then stabilize, demonstrating that when the sentence actions encompass a sufficient range of perspectives for each pattern, the performance of prompt personalization can be sustained at a high level.
@@ 4.7. Case Study
Through the case study presented in Figure 789, and 10, we provide an intuitive demonstration of how variations in these prompts impact the ranking results. For the user in Figure 7, personalizing the “role-playing” pattern from “You are a movie expert” to “ You are a movie ranker to match user’s interest” results in a higher ranking of the ground truth item that the user interacts with next. This illustrates that assigning LLMs as movie rankers to analyze users’ preferences can lead to more precise recommendations for this user. By personalizing the ’history record’ pattern from 5 interaction records to 2 historical records, the ranking results become more precise for the user in Figure 8. This demonstrates that considering a shorter-term interaction history is more effective in capturing his preferences than a long-term one. For the user in Figure 9, personalizing the “reasoning guidance” pattern from “think step by step” to “ update the film preferences before ranking” leads to better recommendations, proving that utilizing LLMs’ refinement and update mechanism is more reasonable than deducing without specific instructions for him. Moreover, personalizing the “output format” pattern from “giving the output format example” to “answering without any unnecessary steps” makes it easier for LLMs to understand the instructions and recommend for the user in Figure 10. We can observe that due to LLMs’ sensitivity to prompts and users’ inner differences, personalizing prompts for each user enables LLMs to generate recommendation results that better satisfy the users, with the ground truth prioritized. Furthermore, we visualize the distribution of users across different patterns under our experiments’ setting to illustrate user diversity vividly, as depicted in Figure 111213, and 14, underscoring the necessity of personalizing each pattern in the prompts for different users. This further validates the significance of instance-wise prompting for recommendation tasks and the effectiveness of our framework.
@@ 4.8. Timing and Computational Complexity Analysis
Compared to directly applying fixed prompts, the streamlined architecture of MARL, with parameters negligible relative to LLMs, incurs minimal additional computational and time costs. When customizing the prompt for each user, compared with heuristical crafting methods like enumeration or token-level optimization approaches like RLPrompt (Deng et al., 2022) and AutoPrompt (Shin et al., 2020), the sentence-level optimization can improve searching efficiency and obtain optimal prompts with minimal iterations, thus reducing timing and computational complexity. However, the multi-round iteration with LLMs is still required since achieving the desired prompt in a single step with MARL is challenging. Formally, let $N$ denote the number of users, $L$ denote the length of a prompt, $D$ denote the possible tokens, and $S$ demote the limited actions. RPP/RPP+ reduces the timing and computational complexity of designing prompts from $O(NLD)$ to $O(NS)$. Our experiments reveal that ChatGPT can complete prompt optimization in approximately three rounds, whereas LLaMa2-7B-chat requires two, seven, and eight iterations on three distinct datasets (i.e., ML-1M, Games, and Lastfm), and Alpaca takes around 8 iterations. The timing and computational complexity are acceptable for achieving significant performance improvements. It reveals that RPP and RPP+ can achieve a balance between prompt quality and resource efficiency.
@@ 5.1. LLMs for Recommender Systems
Recently, large language models (LLMs) have demonstrated remarkable capabilities in instruction following and intent reasoning, drawing researchers’ attention to using LLMs as recommender systems (Wu et al., 2024; Gao et al., 2024; Wu et al., 2022; Mao et al., 2025) which can capture user preference and recommend for them. Conceptualizing LLMs as recommender systems to generate recommendations relies heavily on the quality of the prompts designed. Tailoring prompts of LLMs for the recommendation tasks has been explored by recent research. InstructRec (Zhang et al., 2023b) develops a general instruction format that contains users’ preferences, intentions, task requirements, and contextual information. Chat-REC (Gao et al., 2023) recasts user profiles and prior interactions into prompts, enhancing the interactivity and explainability of the recommendation process. LLM4RS (Dai et al., 2023) designs domain-specific prompt format to analyze ChatGPT’s recommendation ability on all three ranking policies, including point-wise, pair-wise, and list-wise ranking. Instead of designing a common prompt format for the specific recommendation task, we personalize instance-wise prompt for each user to explore users’ different preferences. Though existing studies (Yu et al., 2024; Wang et al., 2024a; Sun et al., 2024) have explored optimizing prompts for users, they may be limited in their reliance on LLMs to refine or partial optimization from limited perspectives. For example, Peprec (Yu et al., 2024) refines the initial prompts by incorporating interactions from similar users in the same cluster iteratively. Re2LLM (Wang et al., 2024a) and ISR (Sun et al., 2024) correct LLMs’ recommendation errors through self-reflection and reasoning. Different from them, our method decomposes the completed prompt into multiple key patterns and optimizes each pattern comprehensively with reinforcement learning.
@@ 5.2. Prompt Sensitivity of LLMs
Recent studies (Lu et al., 2022) have examined the sensitivity of LLMs to prompts that even minor variations in prompts can lead to significantly different outputs, indicating that the one-size-fits-all prompt may not generalize well (Liu et al., 2023a; Mao et al., 2024; Sui et al., 2024) for different inputs. For example, KATE (Liu et al., 2022b) confirmed that the number and order of in-context examples can influence the output of LLMs. Besides, variations in prompt formats within task descriptions may result in LLMs interpreting input instances differently, consequently influencing the outcomes (Zhao et al., 2021). These studies underscore the importance of optimizing prompts for specific inputs to enhance the performance of LLMs. Fixed prompt templates that assume uniformity across all samples in a task, regardless of their varying difficulty, may impair LLMs’ abilities due to a lack of instance-specific knowledge (Jiang et al., 2022). To address the difference in inputs, recent research has proposed instance-level prompt optimization (Jin et al., 2023), which involves rewriting the prompt for each input to better leverage the LLM’s capabilities for specific instances. IPL (Jin et al., 2023) assumed that each learnable prompt token contributes differently to various instances, which can be learned by calculating relevance scores. To meet the specific demands for personalization in recommendation tasks, we enhance LLMs’ personalized recommendations by leveraging instance-wise prompts and introducing a method to optimize these prompts using MARL.
@@ 5.3. Discrete Prompt Engineering
Several approaches have explored prompt engineering in continuous and discrete space to obtain better answers from LLMs. Since the prompts for recent LLMs (ChatGPT, LLaMa2) are discrete and hard prompts, our work is strongly related to the discrete prompts engineering. Prior works have attempted to construct discrete prompts manually (Petroni et al., 2019) or by heuristic (Jiang et al., 2020), which are limited by human effort or resource consumption. Recent research has explored constructing prompts automatically. For instance, prompt retrieval (Rubin et al., 2022) leverages annotated data and a language model, training a compact retriever to retrieve prompts based on input-output pairs. AutoPrompt (Shin et al., 2020) utilizes gradient-guided search to identify the optimal tokens within the prompt, although these prompts are typically not human-interpretable. RLPrompt (Deng et al., 2022) introduces a framework based on RL to generate prompts iteratively word-by-word from the vast vocabulary. An important distinction between our method and the existing methods is that we decompose the completed prompt into multiple key patterns, optimizing each pattern and concatenating them, which can enhance the efficiency of optimization and ensure the quality of prompts. To the best of our knowledge, our work is the first to explore instance-wise prompting in the domain of RSs.
@@ 7. Conclusion
In this paper, we propose instance-wise prompting to personalize prompts for individual users in recommendation tasks. To solve this task, we introduce a new framework, Reinforced Prompt Personalization (RPP/RPP+), personalizing essential patterns respectively with multi-agent and then concatenating them into a personalized prompt. We optimize the prompts at the sentence level and carefully design the expressions for each pattern, enhancing search efficiency and ensuring prompt quality. Experimental results validate the superiority of RPP/RPP+ over several traditional recommender models, few-shot methods, and prompt-based methods in ranking tasks, demonstrating the potential of personalizing instance-wise prompts.
8. Acknowledgements
This research is supported by the National Natural Science Foundation of China (No.92270114, No.62302321, No.U24B20180) and the advanced computing resources provided by the Supercomputing Center of the USTC.